Mathematics 134 – Intensive Calculus for Science 2 Quiz 2 – February 3, 2006

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes. No calculators on this quiz.

A) (10) The following graph shows y = f(x). Use it to sketch the graph y = F (x), where F is an antiderivative of f satisfying F (0) = 1.

B) (7.5) Compute ∫ x^{6} − 4\sqrt{x} + \frac{2}{x} dx.

C) (7.5) Find the solution of \frac{dy}{dx} = cos(x) + 3e^{x} with y(0) = 2.